# Multi-Agent Causal Discovery Using Large Language Models

### v1

@@ 4.3 Results
For the DWD climate data, Meta Agents Model outperforms other methods with the lowest SHD of 5, NHD of 0.194, FNR of 0.833, FPR of 0.2, Precision of 0.333, Recall of 0.166, and F1 score of 0.222. This indicates strong structural accuracy and a balanced performance in terms of precision and recall. The traditional methods such as PC and DirectLiNGAM have higher SHD values (9 and 10, respectively) and lower overall performance in other metrics. The Single-agent zero-shot prompting methods, including GPT-4o, show higher SHD values (10 and 11, respectively) and less competitive performance across the other metrics.
For Sachs’ protein data, Single-agent zero-shot prompting (GPT-4o) stands out with an SHD of 18, the lowest among all methods, indicating the highest structural accuracy. This method also excels in the False Negative Rate (FNR), achieving the best result of 0.098, which signifies a minimal number of missed detections. Its FPR is 0.214, placing it in a competitive position among the top-performing methods. However, while its Precision is 0.230 and F1 score is 0.187, these values reflect a moderate balance between precision and recall. For the GPT-3.5-turbo method, the Coding-Debating Hybrid Group demonstrates notable performance with an SHD of 23, the second-best result in the table. This method also achieves the best FPR of 0.198 and an impressive FNR of 0.068, highlighting its effectiveness in minimizing both false positives and false negatives. The precision of this group is 0.222, and the F1 score is 0.142, indicating a robust overall performance.
In summary, classical stastitcal methods like the PC algorithm, Exact Search, and DirectLiNGAM exhibit higher SHD values, ranging from 24 to 31, and varied performance across FPR, FNR, Precision, and F1 score. Methods integrating LLMs with KBCI, such as PC LLM-KBCI, ES LLM-KBCI, and DirectLiNGAM LLM-KBCI, show moderate improvements but do not surpass the results achieved by Single-agent zero-shot prompting (GPT-4o). Finally, our proposed multi-agent framework has the best performance.
Model SHD NHD FNR FPR Precision Recall F1
1. Auto MPG data with 5 continuous variables
PC 8 0.48 0.8 0.4 0.111 0.2 0.142
Exact Search 7 0.44 0.8 0.35 0.125 0.2 0.153
DirectLiNGAM 8 0.48 0.8 0.4 0.111 0.2 0.142
PC LLM-KBCI 7 0.44 0.8 0.35 0.125 0.2 0.1538
ES LLM-KBCI 7 0.44 0.8 0.35 0.125 0.2 0.1538
DirectLinGam LLM-KBCI 7 0.4 0.6 0.35 0.222 0.4 0.285
Single-agent zero-shot prompting 7 0.28 0.8 0.15 0.25 0.2 0.222
Single-agent zero-shot prompting(GPT-4o) 8 0.36 0.6 0.3 0.25 0.4 0.307
Coding Agents Model (our) 4 0.48 0 0.45 0.357 1 0.526
Meta Agents Model (our) 5 0.2 0.4 0.15 0.5 0.6 0.545
Meta-Coding Hybrid(our) 8 0.48 0.8 0.4 0.111 0.2 0.142
Coding-Meta Hybrid (our) 6 0.32 0.4 0.3 0.333 0.6 0.428
2. DWD climate data with 6 continuous variables
PC 9 0.305 0.833 0.2 0.142 0.167 0.153
Exact Search 6 0.194 0.167 0.2 0.454 0.833 0.588
DirectLinGAM 10 0.388 0.666 0.333 0.166 0.333 0.222
PC LLM-KBCI 7 0.222 0.666 0.133 0.333 0.333 0.333
ES LLM-KBCI 7 0.222 0.333 0.2 0.4 0.666 0.5
DirectLinGam LLM-KBCI 9 0.305 0.333 0.3 0.3076 0.666 0.421
Single-agent zero-shot prompting 10 0.361 1.0 0.233 0 0 0
Single-agent zero-shot prompting(GPT-4o) 11 0.388 1.0 0.266 0 0 0
Coding Agents Model (our) 9 0.305 0.833 0.2 0.142 0.166 0.153
Meta Agents Model (our) 5 0.194 0.833 0.06 0.333 0.1666 0.222
Meta-Coding Hybrid (our) 7 0.277 0.666 0.166 0.285 0.333 0.307
Coding-Debating Hybrid (our) 7 0.25 0.833 0.133 0.2 0.166 0.182
3. Sachs’ protein data with 11 continuous variables
PC 24 0.206 0473 0.156 0.384 0.526 0.444
Exact Search 31 0.330 0.684 0.264 0.181 0.315 0.230
DirectLiNGAM 29 0.289 0.473 0.254 0.277 0.526 0.363
PC LLM-KBCI 30 0.314 0.736 0.225 0.178 0.263 0.212
ES LLM-KBCI 31 0.33 0.68 0.264 0.181 0.315 0.23
DirectLinGam LLM-KBCI 29 0.289 0.473 0.254 0.277 0.5263 0.3636
Single-agent zero-shot prompting 31 0.363 0.526 0.333 0.209 0.473 0.29
Single-agent zero-shot prompting(GPT-4o) 18 0.214 0.842 0.098 0.230 0.157 0.187
Coding Agents Model (our) 29 0.28 0.47 0.25 0.277 0.526 0.363
Meta Agents Model (our) 35 0.33 0.842 0.235 0.111 0.157 0.13
Meta-Coding Hybrid (our) 28 0.272 0.421 0.245 0.305 0.587 0.4
Coding-Meta Hybrid (our) 23 0.198 0.894 0.068 0.222 0.105 0.142
5 Conclusion and Future Research
@@
In this study, a novel framework MAC that integrates agentic workflows of large language models (LLMs) with data-driven methods is introduced. To the best of our knowledge, this is the first work to explore these agentic workflows of LLMs in a causal context. Our framework enhances causal discovery by combining the capabilities of LLMs with empirical data analysis. We acknowledge the need for further research to explore different cooperative workflows, various prompting techniques, and applications in domains such as healthcare, economics, and social sciences. Additionally, recognizing the potential bias in LLMs, more experiments with diverse datasets are required to verify our findings. Preliminary results demonstrate the potential of our framework to improve the accuracy and interpretability of causal discovery. We hope that our work lays a foundational stone for future research, inspiring advancements in the integration of LLMs with causal inference methodologies and contributing to more informed decision-making and policy development.
A.1 Meta-Debate Module Visualization
@@
A.2 Debate-Coding Module Visualization
A.2.1 Phase 1: Plan Debating
@@
A.2.2 Phase 2: Code Executing
@@
Custom Appendix Title B Details prompting design of each agent in MAC
We use AutoGen (Wu et al. [2023]) framework to implement our methods
B.1 Meta-Debate Module
B.1.1 Debaters’ Prompt
@@ B.1.2 Judge Prompt
B.2 Debate-Coding Module
B.2.1 Plan Debaters Prompt
@@ B.2.2 Judge Prompt
B.2.3 Code Executor Prompt
@@
Custom Appendix Title C Sample Answers from Experiments
C.1 Meta Agents Model
@@ 4. Are there any direct causal relationships from Displacement to Mpg(miles per gallon)?
C.2 Coding Agents Model
C.2.1 Phase 1: Plan Debating
@@
C.2.2 Phase 2: Code Executing

### v2

@@ 4.2 Performance of MAC
DWD Climate Dataset: In the DWD climate dataset, which comprises a 6x6 matrix graph, the Causal Agent Debate method again showed strong performance, with GPT-3.5 achieving the lowest SHD of 5 and NHD of 0.194. The Coding Agent method performed well with GPT-4o mini, achieving an SHD of 6 and NHD of 0.194. The Coding-Debating Hybrid method stood out, with Llama-3.1-8b achieving the lowest SHD (5) and NHD (0.138). Similarly, the Hybrid-Debating-Coding Hybrid method also performed well, with GPT-4o achieving an SHD of 5 and NHD of 0.138. The traditional methods such as PC and DirectLiNGAM have higher SHD values (9 and 10, respectively) and lower overall performance in other metrics. The Single-agent zero-shot prompting methods, including GPT-4o, show higher SHD values (10 and 11, respectively) and less competitive performance across the other metrics.
Sachs’ Protein Dataset : In the Sachs’ protein dataset, the Causal Agent Debate method displayed mixed results, with GPT-4o achieving an SHD of 21 and an NHD of 0.231. The Coding Agent method outperformed others, with Llama-3.1-70b achieving the lowest SHD of 18 and an NHD of 0.157. The Coding-Debating Hybrid method demonstrated consistent performance, with GPT-4o mini reaching an SHD of 21 and GPT-3.5 achieving the lowest NHD of 0.198. The Debating-Coding Hybrid method also performed well, with GPT-4o mini achieving an SHD of 22 and NHD of 0.190.
Model Auto MPG Climate Sachs
SHD NHD SHD NHD SHD NHD
Other methods
PC 8 0.48 9 0.305 24 0.206
Exact Search 7 0.44 6 0.194 31 0.33
LINGAM 8 0.48 10 0.388 29 0.289
PC LLM-KBCI 7 0.44 7 0.222 30 0.314
ES LLM-KBCI 7 0.44 7 0.222 31 0.33
DirectLiGam LLM-KBCI 7 0.4 9 0.305 29 0.289
Single-agent (GPT-4o) 8 0.36 11 0.388 18 0.214
Single-agent (GPT-3.5) 7 0.28 10 0.361 31 0.363
Causal Agent Debate
GPT-4o 5 0.2 9 0.333 35 0.371
GPT-4o mini 4 0.16 11 0.416 35 0.338
GPT-3.5 5 0.2 5 0.194 21 0.231
Llama 3.1 70b 10 0.44 8 0.222 35 0.380
Llama 3.1 8b 7 0.28 9 0.277 35 0.338
Gemini 2 9b 5 0.2 7 0.222 25 0.223
Coding Agents
GPT-4o 8 0.48 9 0.388 30 0.322
GPT-4o mini 6 0.6 6 0.194 30 0.322
GPT-3.5 4 0.48 9 0.305 29 0.28
Llama 3.1 70b 6 0.6 6 0.194 18 0.157
Llama 3.1 8b 6 0.6 9 0.388 36 0.396
Gemini 2 9b (GPT-4o-mini’s plan) 7 0.36 - - - -
Hybrid-Coding-Debating
GPT-4o 6 0.28 10 0.305 23 0.247
GPT-4o mini 5 0.24 9 0.25 21 0.190
GPT-3.5 6 0.32 7 0.25 23 0.198
Llama 3.1 70b 6 0.36 6 0.166 33 0.314
Llama 3.1 8b 5 0.2 5 0.138 33 0.305
Gemini 2 9b - - - - - -
Hybrid-Debating-Coding
GPT-4o 5 0.24 5 0.138 29 0.297
GPT-4o mini 8 0.48 7 0.25 22 0.190
GPT-3.5 8 0.48 7 0.277 28 0.272
Llama 3.1 70b 5 0.24 6 0.194 29 0.289
Llama 3.1 8b 6 0.4 9 0.388 - -
Gemini 2 9b - - - - - -
Table 2: Combined results across Auto MPG, Climate, and Sachs Datasets
4.3 Performance among LLMs
Small size LLMs: Llama 3.1 8b is a strong small model, consistently achieving solid results in hybrid models and competing with larger models in simpler datasets like Auto MPG and Climate. It excels in Hybrid Coding-Debating settings. Gemini 2 9b, while capable, tends to lag behind Llama 3.1 8b, especially since the MAC models require coding, particularly when the prompting token count exceeds 8,000 to 9,000 tokens, which leads to random code generation and prolonged debugging loops. Therefore, we do not include the result of this model in our research.
Medium size LLM: Llama 3.1 70b is a versatile medium-sized model, excelling particularly in coding tasks and holding its own in hybrid models. It tends to perform better in more complex datasets like Sachs, where it even surpasses some larger models. This model strikes a balance between computational efficiency and strong performance.
Large size LLMs: GPT-4o mini  [^2] If GPT-4o mini were considered a small size model, it would be considered as the best model in this setting is an outstanding performer across hybrid and debate models, particularly in simpler datasets like Auto MPG and Climate, where it achieves some of the lowest SHD and NHD scores. GPT-4o is similarly strong in hybrid tasks and coding tasks, making it a top-tier choice for flexible, multi-agent collaboration. GPT-3.5, while effective, tends to be outperformed by the GPT-4 models in hybrid tasks but remains strong in causal debates and coding tasks for simpler datasets.
4.4 Token Usages and LLMs insights on MAC
Figure 4: Tokens Usage Prompting-Completion Tokens Ratio
4.4.1 Debate-Coding Module
Based on overall token consumption (Figure 4), (1) Coding-Debating Hybrid and Debating-Coding Hybrid models consumed the highest number of tokens. (2) The Agents-only model used a moderate number of tokens. The Coding Agents model consistently required the fewest tokens, even for more complex datasets and larger language models.
For smaller datasets like 5x5 and 6x6 matrices, the Agents-only and Coding-Debating Hybrid models used 50,000 to 60,000 tokens. However, with larger datasets, such as the 11x11 matrix, token usage increased significantly to about 230,000 to 300,000 tokens. To reduce this, future implementations can limit debating rounds to one or two, as models occasionally initiated five or six rounds, leading to excessive token use. Conversely, the Coding Agents and Debating-Coding Hybrid models showed more stable token consumption, ranging from 40,000 to 100,000 tokens, with performance influenced by their ability to generate and debug code. While GPT-series models and Llama-3.1 series performed reliably, smaller models like Gemini 2-9b struggled with prompts exceeding 8,000 to 9,000 tokens, leading to random code generation. Surprisingly, Llama-3.1-8b outperformed others in Gemini 2-9b generation, except the last dataset, despite requiring more tokens for debugging. This suggests it can effectively utilize plans from larger models, which typically require fewer tokens. Further experiments will be detailed in the next Ablation Studies 4.5 section.
The prompt-to-completion token ratio varied across models (see Figure 4): the Agents-only and Coding-Debating Hybrid models had a 2:1 ratio (2,000 prompting tokens to 1,000 completion tokens), while the Coding Agents and Debating-Coding Hybrid models had a 10:1 ratio (10,000 prompting tokens to 1,000 completion tokens). This significant gap in the Coding Agents and Debating-Coding Hybrid models arises from the intensive debugging process, where even small codebases can generate large outputs as the system traces the source code, leading to increased token consumption, especially during debugging
4.5 Ablation Study
Single-agent vs Causal Agent Debate: Causal debating agents outperform single-agent models in handling complexity by leveraging multiple viewpoints, making them better suited for more intricate problems. For example, in the Auto MPG dataset, Causal Agent Debate GPT-4o achieves an SHD of 5, outperforming the Single-agent GPT-4o with an SHD of 8. In the Climate dataset, for example, Causal Agent Debate GPT-4o achieves an SHD of 9, while Single-agent GPT-4o achieves 11. However, in complex datasets, such as Sachs, single-agent GPT-3.5 has an SHD of 31 and an NHD of 0.363, while Causal Agent Debate GPT-4o performs worse (SHD = 35), but the debate framework at least provides a structured method to tackle the challenge, despite not being fully optimized here.
Model Auto Dataset Climate Dataset Sachs Dataset
SHD NHD SHD NHD SHD NHD
Single-agent (GPT 3.5) 7 0.28 10 0.361 31 0.363
Single-agent (GPT-4o) 8 0.36 11 0.388 18 0.214
Our single-agent (GPT-4o mini) 6 0.28 9 0.361 24 0.289
Table 3: Comparision between various single-agent designs
Eliminating the Judge: We retained only one causal single agent (GPT-4o-mini) with our curated prompt, as shown in Figure 3. This causal-single-agent did not perform better than other MAC models, except the Agents-only Model in the Sachs dataset. However, compared with other agents such as GPT-3.5 and GPT-4o, which use 0-CoT and ICL techniques, our single-agent model achieved better performance but only stayed behind GPT-4o on the Sachs dataset given the complexity of the dataset.
Model Dataset SHD NHD Condition
Gemini 2 9b Auto MPG 0 0 Baseline
Gemini 2 9b Auto MPG 7 0.36 With GPT-4o-mini’s plan
Llama 3.1 8b Sachs 36 0.396 Baseline
Llama 3.1 8b Sachs 18 0.157 With Llama 3.1 70b’s plan
Table 4: Performance of Gemini 2 9b and Llama 3.1 8b combining superior models
Combining small and large models: In the Coding Agents Model, we experimented with the combination of using the plan from GPT-4o and coding from Gemini 2-9b. With only an exception in the dataset Auto MPG, the model was able to complete and implement the plan successfully but was not able to precede the rest of the other datasets. However, as for the Llama-3.1-70b and Llama-3.1-8b, it exhibits promising performance. Specifically, after observing that Llama-3.1-70b achieved the best performance with its proposed plan in the Sachs dataset, we used that plan to prompt Llama-3.1-8b to implement it. Surprisingly, the smaller model yielded similar performance. Therefore, if computational resources are limited, it is advisable to use larger models for high-level tasks such as planning and smaller models (Coding Agents) for computationally intensive tasks like coding and debugging.
5 Conclusion
@@
In this study, we introduce a novel framework, MAC, that integrates the agentic workflows of large language models (LLMs) with data-driven methods. To the best of our knowledge, this is the first investigation into the agentic workflows of LLMs within a causal context. Our framework enhances causal discovery by synergizing the capabilities of LLMs with empirical data analysis. We propose three distinct models that leverage the causal reasoning abilities of LLMs alongside observational data. Additionally, we conducted extensive experiments across various LLM sizes, analyzing token consumption and developing strategies to address related challenges. We recognize the necessity for further research to explore across domains such as healthcare, economics, and social sciences. We hope our work serves as a foundational stone for future research, inspiring advancements in the integration of LLMs with causal inference methodologies and contributing to more informed decision-making and policy development.
Appendix A Appendix
A.1 Additionally experiment results
A.1.1 Auto MPG data
Auto MPG data (Quinlan (1993)): This dataset consists of the variables around the fuel consumption of cars. With five variables: “Weight”, “Displacement”, “Horsepower”, “Acceleration” and “Mpg”(miles per gallon)
Model SHD NHD
Other methods
PC 8 0.48
Exact Search 7 0.44
LINGAM 8 0.48
PC LLM-KBCI 7 0.44
ES LLM-KBCI 7 0.44
DirectLiGam LLM-KBCI 7 0.4
Single-agent (GPT-4o) 8 0.36
Single-agent (GPT-3.5) 7 0.28
Causal Agent Debate
GPT-4o 5 0.2
GPT-4o mini 4 0.16
GPT-3.5 5 0.2
Llama 3.1 70b 10 0.44
Llama 3.1 8b 7 0.28
Gemini 2 9b 5 0.2
Coding Agents
GPT-4o 8 0.48
GPT-4o mini 6 0.6
GPT-3.5 4 0.48
Llama 3.1 70b 6 0.6
Llama 3.1 8b 6 0.6
Gemini 2 9b (gpt-4o-mini’s plan) 7 0.36
Hybrid-Coding-Debating
GPT-4o 6 0.28
GPT-4o mini 5 0.24
GPT-3.5 6 0.32
Llama 3.1 70b 6 0.36
Llama 3.1 8b 5 0.2
Gemini 2 9b - -
Hybrid-Debating-Coding
GPT-4o 5 0.24
GPT-4o mini 8 0.48
GPT-3.5 8 0.48
Llama 3.1 70b 5 0.24
Llama 3.1 8b 6 0.4
Gemini 2 9b - -
Table 5: Results on Auto MPG Dataset
A.1.2 DWD climate data
DWD climate data (Mooij et al. (2016)): This dataset encompasses six continuous variables capturing climate observations such as altitude, temperature, precipitation levels, longitude, sunshine duration, and latitude. It is aimed at studying weather patterns, climate change impacts, and geographical correlations in climate variables.
Model SHD NHD
Other methods
PC 9 0.305
Exact Search 6 0.194
LINGAM 10 0.388
PC LLM-KBCI 7 0.222
ES LLM-KBCI 7 0.222
DirectLiGam LLM-KBCI 9 0.305
Single-agent (GPT-4o) 11 0.388
Single-agent (GPT-3.5) 10 0.361
Causal Agent Debate
GPT-4o 9 0.333
GPT-4o mini 11 0.416
GPT-3.5 5 0.194
Llama 3.1 70b 8 0.222
Llama 3.1 8b 9 0.277
Gemini 2 9b 7 0.222
Coding Agents
GPT-4o 9 0.388
GPT-4o mini 6 0.194
GPT-3.5 9 0.305
Llama 3.1 70b 6 0.194
Llama 3.1 8b 9 0.388
Gemini 2 9b - -
Hybrid-Coding-Debating
GPT-4o 10 0.305
GPT-4o-mini 9 0.25
GPT-3.5 7 0.25
Llama 3.1 70b 6 0.166
Llama 3.1 8b 5 0.138
Gemini 2 9b - -
Hybrid-Debating-Coding
GPT-3.5 7 0.277
GPT-4o 5 0.138
GPT-4o mini 7 0.25
Llama 3.1 70b 6 0.194
Llama 3.1 8b 9 0.388
Gemini 2 9b - -
Table 6: Results on Climate Dataset
A.1.3 Sachs protein data
Sachs protein data (Sachs et al. (2005)): The dataset comprises protein signaling measurements from multiparameter single-cell data, capturing the interactions among various proteins (raf, mek, plc, pip2, pip3, erk, akt, pka, pkc, p38, jnk). It’s aimed at understanding signal transduction pathways within cells, derived from an influential study published in Science.
Model SHD NHD
Other methods
PC 24 0.206
Exact Search 31 0.33
LINGAM 29 0.289
PC LLM-KBCI 30 0.314
ES LLM-KBCI 31 0.33
DirectLiGam LLM-KBCI 29 0.289
Single-agent (GPT-4o) 18 0.214
Single-agent (GPT-3.5) 31 0.363
Causal Agent Debate
GPT-4o 35 0.371
GPT-4o mini 35 0.338
GPT-3.5 21 0.231
Llama 3.1 70b 35 0.380
Llama 3.1 8b 35 0.338
Gemini 2 9b 25 0.223
Coding Agents
GPT-4o 30 0.322
GPT-4o mini 30 0.322
GPT-3.5 29 0.28
Llama 3.1 70b 18 0.157
Llama 3.1 8b 36 0.396
Gemini 2 9b - -
Hybrid-Coding-Debating
GPT-4o 23 0.247
GPT-4o-mini 21 0.190
GPT-3.5 23 0.198
Llama 3.1 70b 33 0.314
Llama 3.1 8b 33 0.305
Gemini 2 9b - -
Hybrid-Debating-Coding
GPT-4o 29 0.297
GPT-4o mini 22 0.190
GPT-3.5 28 0.272
Llama 3.1 70b 29 0.289
Llama 3.1 8b - -
Gemini 2 9b - -
Table 7: Results on Sachs Dataset
A.2 Visualization of MAC modules
A.2.1 Meta-Debate Module Visualization
@@
A.2.2 Debate-Coding Module Visualization
Phase 1: Plan Debating
@@ A.2.2 Debate-Coding Module Visualization
Phase 2: Code Executing
@@
A.3 Details prompting design of each agent in MAC
We use AutoGen (Wu et al. (2023)) framework to implement our methods
A.3.1 Meta-Debate Module
@@
A.3.2 Debate-Coding Module
Plan Debaters Prompt
@@ A.3.2 Debate-Coding Module
Code Executor Prompt
@@
Appendix B Sample Answers from Experiments
B.0.1 Meta Agents Model
@@ 4. Are there any direct causal relationships from Displacement to Mpg(miles per gallon)?
B.0.2 Coding Agents Model
Phase 1: Plan Debating
@@ B.0.2 Coding Agents Model
Phase 2: Code Executing